OpenStudy (anonymous):

A dehydrated patient needs a 3.6% saline IV. Unfortunately, the hospital only has bags of 2% and 4% saline solutions. How many liters of each of these solutions should be mixed together to yield 1 liter of the desired concentration?

OpenStudy (anonymous):

\[.04x+.02(1-x)=.036\] should do it

hero (hero):

Whatever happened to using the beaker method?

hero (hero):

I showed it to you once

OpenStudy (anonymous):

that was wrong \[40x+20(1-x)=36\] is what i meant

OpenStudy (anonymous):

my short term memory got shot off in the war

hero (hero):

lol, I guess

OpenStudy (anonymous):

not saying which war though... @shesmketr you got it from here?

OpenStudy (anonymous):

What he said...

OpenStudy (anonymous):

\[40x+20-20x=36\] \[20x+20=36\] \[20x=16\] \[x=\frac{16}{20}=.8\] so .8 of the 4% solution, .2 of the 2%

OpenStudy (anonymous):

and no 7% solution